DK4L:

Which statement concerning the equation x^2+1=2x is true? Its discriminant is -8, so its solutions are negative. Its discriminant is 0, so it has one real solution. Its discriminant is 0, so it has no solution. Its discriminant is -8, so it has two complex solutions.

smackz:

Rewrite the equation: -2x^2 - 3x + 8 = 0 2x^2 + 3x -8 =0 Where a=2, b=3 and c=-8 Then b^2 - 4ac = 3^2 - 4(2)(-8) = 9 + 64 = 73 A positive discriminant implies that the equation has two different real solutions. Answer: the discriminant is 73, so the equation has 2 real solution

DK4L:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @smackz Rewrite the equation: -2x^2 - 3x + 8 = 0 2x^2 + 3x -8 =0 Where a=2, b=3 and c=-8 Then b^2 - 4ac = 3^2 - 4(2)(-8) = 9 + 64 = 73 A positive discriminant implies that the equation has two different real solutions. Answer: the discriminant is 73, so the equation has 2 real solution \(\color{#0cbb34}{\text{End of Quote}}\) ok thnx

jigglypuff314:

Please realize that the equation smackz has used is not the equation that DK4L asked for but the theory / ideas / steps are the same

jigglypuff314:

DK4L, "2 real solutions" is NOT your answer btw

DK4L:

oh

jigglypuff314:

smackz used the equation: 2x^2 + 3x -8 =0 which is not what you wanted

DK4L:

oh so what is the answer?

jigglypuff314:

x^2+1=2x can be rearranged to look like: x^2 - 2x + 1 = 0 This matches up with the "standard form" of quadratic equations: ax^2 + bx + c = 0 Which means that a=1 , b=-2, and c=1 What do you get if you plug in a, b, and c into the formula for the discriminant? b^2 - 4ac

DK4L:

yeahhhhhh ok i get it now

jigglypuff314:

Ok, if you would like for me to double check your answer, just let me know

DK4L:

Oh K

DK4L:

ight i got it thanks

darkknight:

@dude move to correct subject please